OpenStudy (vishweshshrimali5):

The solution set of \(\large{(2\sqrt{2} - 2)^{x} + (6-4\sqrt{2})^{x} \ge 2^{1+x}}\) is ??

OpenStudy (vishweshshrimali5):

@iambatman @ganeshie8 @dan815

mathslover (mathslover):

You always forget one user to tag ... @mathslover

mathslover (mathslover):

This is easy.

mathslover (mathslover):

First step is clear, cancel 2^x from both sides.

mathslover (mathslover):

\((\sqrt{2} -1)^x + (3-2\sqrt{2})^x \ge 2\)

OpenStudy (vishweshshrimali5):

You got the first step right !!

mathslover (mathslover):

Now... use some logic. Clearly : \(\sqrt{2} - 1 < 1\) | and \((3-2\sqrt{2}) < 1\)

OpenStudy (vishweshshrimali5):

Soo ?

mathslover (mathslover):

So, : for x not to be negative : \((\sqrt{2} - 1)^x < 1 \) and \((3-2\sqrt{2})^x < 1\) Add these two equations, you get what you are not given. So, x is not positive. Now, it should be clear that x is negative.

OpenStudy (vishweshshrimali5):

You are wrong ;)

mathslover (mathslover):

Hmm , wait, let me recheck what I did.

mathslover (mathslover):

Are you pointing towards the condition of x = 0 ?

OpenStudy (vishweshshrimali5):

nope

mathslover (mathslover):

x is negative or x = 0 ...

mathslover (mathslover):

And I know, that I'm not wrong here.

OpenStudy (vishweshshrimali5):

Nope

OpenStudy (vishweshshrimali5):

Yeah sorry you are right.

mathslover (mathslover):

Okay, so, we can say that x is negative or x = 0

mathslover (mathslover):

Now, to find the solution set, we will have to do something creative.